OpenStudy (anonymous):

What is the solution of the equation 4x + 17 = 45?

OpenStudy (accidentalaichan):

4x + 17 = 45 Reorder the terms: 17 + 4x = 45 Solving 17 + 4x = 45 Move all terms containing x to the left, all other terms to the right. Add '-17' to each side of the equation. 17 + -17 + 4x = 45 + -17 Combine like terms: 17 + -17 = 0 0 + 4x = 45 + -17 4x = 45 + -17 Combine like terms: 45 + -17 = 28 4x = 28 Divide each side by '4'. x = 7 Simplifying x = 7